# MTF-Grasp: A Multi-tier Federated Learning Approach for Robotic Grasping

### v1

@@ I INTRODUCTION
Robotic grasping focuses on manipulators grasping objects effectively, and has found potential application to mainstream robotic tasks, including warehouse automation, search and rescue, as well as manufacturing and assembly [1]. Increasingly, robotic tasks necessitate a group of robots to perform similar tasks across various aforementioned applications. Thus, the need for collaboration in robotic fleets is evident [2]. However, most current robotic grasping algorithms focus on stand-alone and cloud-based centralized training scenarios. In the former scenario, the robots individually train models with their local datasets, which can be used locally for inference [3]. While this scenario removes the need for communication and maintains privacy, it does not benefit collaboration and knowledge sharing among robots. Conversely, cloud-based training involves robots transferring raw data to the cloud, where a centralized model is trained. Inference in cloud-based grasping can be carried out by either the cloud or the models can be transferred to robots for local inference [4]. While providing collaboration, cloud-based training risks privacy [5] and increases communication load, as it requires data transfer to the cloud. Federated learning (FL) [6] emerges as a promising middle ground, enabling collaboration while preserving data privacy.
@@
While FL enables collaboration between devices, several challenges persist, including data and system heterogeneity, communication cost, and privacy [9]. Data heterogeneity is one of the main challenges in the robotic domain [10], as different robots may own i) data from different distributions or ii) different quantities of data, both cases representing non-independent and identically distributed (non-IID) data [11]. Quantity-skewed data, where the amount of data across different devices is imbalanced, can lead to inaccuracy and limited generalization in FL [12]. Furthermore, data quantity skewness introduces clients with limited data. Limited data can lead to local model overfitting, which is crucial in robotic manipulation tasks, as collecting new data is costly and data augmentation is challenging due to the complex nature of these tasks. To mitigate problems arising from non-IID and, specifically, quantity data skewness in robotic grasping, we propose MTF-Grasp. This multi-tier FL approach serves as a knowledge transfer mechanism between robots, reducing model overfitting in robots with limited data. MTF-Grasp ranks robots based on data quality and quantity, dividing them into top-level and low-level robots. It utilizes the better data distribution in top-level robots to improve models in low-level robots. The summary of our contributions is as follows:
@@ IV SYSTEM DESIGN
Consider a set $\mathcal{R}$ of $n$ robots, $\mathcal{R}$, where each robot collects and learns from its local dataset $\mathcal{D}^{r\in\mathcal{R}}$, and the cumulative dataset can be represented as $\mathcal{D}=\sum_{r\in\mathcal{R}}{\mathcal{\mathcal{D}}^{r}}$. The total set of classes or categories in the dataset $\mathcal{D}$ is $\mathcal{C}$ of size $m$, and $\mathcal{D}^{r}_{c}$ represents the data of class $c\in\mathcal{C}$ belonging to robot $r\in\mathcal{R}$. We assume that the data distribution is skewed among the robots. To tackle the non-IID nature of data, specifically data quantity skewness, MTF-Grasp selects a subset of robots for initial training based on data quality and quantity distribution. The client selection and training processes are discussed below and depicted in Figure 2.
@@ IV-A1 Data Quality
From a data quality perspective, the robots with the lowest deviation amongst the samples of different global classes are considered higher-quality robots. Consider two robots who own (3,100,2), and (52,76,88) samples for three different global classes respectively. As the deviation among number of samples for the second robot is lower, the data quality in that robot is deemed to be higher. The data distribution score that quantifies quality is calculated as Eqn. 4 for each robot.
@@ V-A Datasets
In this work, we use the Cornell [29] and Jacquard [30] grasping datasets, two of the most widely used robotic grasping datasets. Both datasets contain RGB-D images alongside annotations for the grasping areas in each image. The Cornell dataset contains 885 RGB-D images of around 240 different real objects. The Jacquard dataset has around 54,000 images from around 11,500 objects. The Jacquard dataset contains images from ShapeNet [31], which is a large CAD model dataset. For our experiments, we use a subset of the Jacquard datasets that contains around 4000 distinct objects, hereafter denoted as Jacquard-4k.
To simulate the non-IID distribution of data, the dataset must be divided into classes, and while each object category can act as a class, this distribution would significantly dilute objects per class. Hence, we divided the Cornell dataset into seven classes for the experiment’s purpose. The choice of the number of classes was inspired by  [28], but we added an extra class for objects that didn’t fit any of the other classes. Table I contains the classes and number of samples for each class. Fig. 3 provides the sample distribution across clients in both class-based non-IID and quantity-skew scenarios for the Cornell dataset.
@@ V-B Model Training and Hyperparameters
This work utilizes the grasping network defined in [32] and the implementation provided by the authors [^1] https://github.com/ZXP-S-works/SE2-equivariant-grasp-learning for the main grasping network with small adjustments and a horizontal FL setup. The experiments were carried out with $n=7$ robots, with a number of top-level robots $j=2$.
Data is divided into local training and test sets for each robot, consisting of 85% and 15% respectively. Accuracy is the ratio of positive grasp predictions from the test set and the number of samples in the test set. MTF-Grasp can be utilized with any aggregation function, and in this case, MTF-Grasp is coupled with FedAvg and FedNova, shown as MTF-Grasp-Avg and MTF-Grasp-Nova in the experiments. For the algorithmic hyperparameter values, we have used $e_{t}=5$ for a number of local iterations in top-level robots, $e_{r}=15$ for the number of iterations in low-level robots, while the number of global communication rounds for FL training is set to 10. Furthermore, we found that the value $0.5$ worked best for both $\lambda_{DDS}$ and $\lambda_{DQS}$.

### v2

@@ I INTRODUCTION
Robotic grasping focuses on manipulators grasping objects effectively and has found potential applications in mainstream robotic tasks [1], including warehouse automation, search and rescue, as well as manufacturing and assembly [2]. Increasingly, robotic tasks necessitate a group of robots to perform similar tasks across various aforementioned applications. Thus, the need for collaboration in robotic fleets is evident [3]. However, most current robotic grasping algorithms focus on stand-alone and cloud-based centralized training scenarios. In the former scenario, the robots individually train models with their local datasets, which can be used locally for inference [4]. While this scenario removes the need for communication and maintains privacy, it does not benefit collaboration and knowledge sharing among robots. Conversely, cloud-based training involves robots transferring raw data to the cloud, where a centralized model is trained. Inference in cloud-based grasping can be carried out by either the cloud or the models can be transferred to robots for local inference [5]. While providing collaboration, cloud-based training risks privacy [6] and increases communication load, as it requires data transfer to the cloud. Federated learning (FL) [7] emerges as a promising middle ground, enabling collaboration while preserving data privacy.
@@
While FL enables collaboration between devices, several challenges persist, including data and system heterogeneity, communication cost, and privacy [10]. Data heterogeneity is one of the main challenges in the robotic domain [11], as different robots may own i) data from different distributions or ii) different quantities of data, both cases representing non-independent and identically distributed (non-IID) data [12]. Quantity-skewed data, where the amount of data across different devices is imbalanced, can lead to inaccuracy and limited generalization in FL [13]. Furthermore, data quantity skewness introduces clients with limited data. Limited data can lead to local model overfitting, which is crucial in robotic manipulation tasks, as collecting new data is costly and data augmentation is challenging due to the complex nature of these tasks. To mitigate problems arising from non-IID and, specifically, quantity data skewness in robotic grasping, we propose MTF-Grasp. This multi-tier FL approach serves as a knowledge transfer mechanism between robots, reducing model overfitting in robots with limited data. MTF-Grasp ranks robots based on data quality and quantity, dividing them into top-level and low-level robots. It leverages the improved data distribution in top-level robots to enhance models in low-level robots. The summary of our contributions is as follows:
@@ IV SYSTEM DESIGN
Consider a set $\mathcal{R}$ of $n$ robots, $\mathcal{R}$, where each robot collects and learns from its local dataset $\mathcal{D}^{r\in\mathcal{R}}$, and the cumulative dataset can be represented as $\mathcal{D}=\sum_{r\in\mathcal{R}}{\mathcal{\mathcal{D}}^{r}}$. The total set of classes or categories in the dataset $\mathcal{D}$ is $\mathcal{C}$ of size $m$, and $\mathcal{D}^{r}_{c}$ represents the data of class $c\in\mathcal{C}$ belonging to robot $r\in\mathcal{R}$. We assume that the data distribution is skewed among the robots. To address the non-IID nature of the data, specifically data quantity skewness, MTF-Grasp selects a subset of robots for initial training based on the distribution of data quality and quantity. The client selection and training processes are discussed below and depicted in Figure 2.
@@ IV-A1 Data Quality
From a data quality perspective, the robots with the lowest deviation amongst the samples of different global classes are considered higher-quality robots. Consider two robots who own (3,100,2), and (52,76,88) samples for three different global classes respectively. As the deviation among the number of samples for the second robot is lower, the data quality in that robot is deemed to be higher. The data distribution score that quantifies quality is calculated as Eqn. 4 for each robot.
@@ V-A Datasets
In this work, we use the Cornell [30] and Jacquard [31] grasping datasets, two of the most widely used robotic grasping datasets. Both datasets contain RGB-D images alongside annotations for the grasping areas in each image. The Cornell dataset comprises 885 RGB-D images of approximately 240 distinct real-world objects. The Jacquard dataset has around 54,000 images from around 11,500 objects. The Jacquard dataset comprises images from ShapeNet [32], a large CAD model dataset. For our experiments, we use a subset of the Jacquard datasets, which contains approximately 4000 distinct objects, hereafter denoted as Jacquard-4k.
To simulate the non-IID distribution of data, the dataset must be divided into classes, and while each object category can act as a class, this distribution would significantly dilute objects per class. Hence, we split the Cornell dataset into seven classes for the experiment’s purpose. The choice of the number of classes was inspired by  [29], but we added an extra class for objects that didn’t fit any of the other classes. Table I contains the classes and number of samples for each class. Fig. 3 provides the sample distribution across clients in both class-based non-IID and quantity-skew scenarios for the Cornell dataset.
@@ V-B Model Training and Hyperparameters
This work utilizes the grasping network defined in [33] and the implementation provided by the authors [^1] https://github.com/ZXP-S-works/SE2-equivariant-grasp-learning for the main grasping network with minor adjustments and a horizontal FL setup. The experiments were conducted with $n=7$ robots, including a subset of top-level robots $j=2$.
Accuracy is the ratio of positive grasp predictions from the test set and the number of samples in the test set. MTF-Grasp can be utilized with any aggregation function, and in this case, MTF-Grasp is coupled with FedAvg and FedNova, shown as MTF-Grasp-Avg and MTF-Grasp-Nova in the experiments. For the algorithmic hyperparameter values, we have used $e_{t}=5$ for a number of local iterations in top-level robots, $e_{r}=15$ for the number of iterations in low-level robots, while the number of global communication rounds for FL training is set to 10. Furthermore, we found that the value $0.5$ worked best for both $\lambda_{DDS}$ and $\lambda_{DQS}$.